OpenStudy (anonymous):

5 over 22 subtracted by 5 over 33

OpenStudy (hoblos):

5/22 - 5/33 = 15/66 - 10/66 = 5/66

OpenStudy (anonymous):

Thank you :)

OpenStudy (hoblos):

welcome :D